# TripleMixer: A 3D Point Cloud Denoising Model for Adverse Weather

### v1

@@ III-A Adverse Weather Simulator
III-B Dataset Statistics
Adverse weather primarily causes back-scattering and attenuation of LiDAR pulse transmissions, leading to significant shifts in both range and intensity for points in a LiDAR point cloud[43, 47]. Figure 1 illustrates the semantic distributions of the same point cloud frame under three different weather conditions within our proposed Weather-KITTI dataset and their corresponding clean datasets. The weather noise points induced by the three weather conditions are primarily concentrated around the LiDAR sensor. Snowflakes contain ice crystals with varying densities, and in snowy conditions, each laser beam interacts with airborne particles, reflecting at different angles and considering potential occlusions. Compared to rain and fog, which consist of denser small water droplets, snowflakes result in more dispersed noise points.
Table II delineates the severity levels of various weather-induced corruptions as configured in this study. In the snow simulation, the snowfall rate $r_{s}$ is set at [0.5, 1.0], [1.5, 2.0], and [2.5, 3.0] mm/h to represent light, moderate, and heavy snowfall, respectively. For fog simulations, the fog parameter $\beta$ is adjusted to [0.01, 0.05], [0.08, 0.14], and [0.18, 0.25] to reflect light, moderate, and heavy fog conditions. Similarly, in the rain simulation, the rain rate $r_{r}$ is set to [1.0, 1.5], [1.8, 2.4], and [2.6, 3.0] mm/h to simulate light, moderate, and heavy rainfall. As the severity of the weather increases, the extent of point cloud corruption proportionally intensifies. According to Table II, all LiDAR frames of Weather-KITTI and Weather-NuScenes are randomly synthesized with three different levels of weather severity in this study. This configuration allows us to investigate the generalizability of during models training. Simultaneously, we ensure that each level of weather severity has a similar number of LiDAR frames, guaranteeing the diversity of data across different weather scenarios. As depicted in Figure 2(a), the number of LiDAR frames for each weather category under the three weather severity levels in Weather-KITTI is approximately 14K frames. Similarly, as shown in Figure 2(b), the number of LiDAR frames for each weather category under each weather severity level in Weather-NuScenes is around 9K frames.
@@ III-B Dataset Statistics
Figure 4: Overview of the proposed denoising model, the TripleMixer network. The TripleMixer network comprises three main layers: the Geometry Mixer (GMX) Layer, the Frequency Mixer (FMX) Layer, and the Channel Mixer (CMX) Layer. These layers are detailed in Section IV.
Weather-KITTI is a large-scale synthetic LiDAR dataset designed for adverse weather conditions, based on SemanticKITTI [38]. It includes 22 sequences with a total of 130,656 LiDAR scans. Each point cloud is segmented into 21 semantic classes, with the semantic classes for snow, fog, and rain mapped to labels 110, 111, and 112, respectively, across all sequences. As SemanticKITTI provides semantic labels only for the first ten sequences, we also limit the availability of other semantic labels to the first ten sequences. Weather-NuScenes, derived from nuScenes-lidarseg [39], contains 84,390 LiDAR scans evenly divided into ten sequences. It includes 19 semantic classes and has similarly been enhanced with semantic classes for snow, fog, and rain across all sequences, mapped to labels 110, 111, and 112, respectively. Figure 3 illustrates the distribution of LiDAR points across semantic categories for all sequences in both Weather-KITTI and Weather-NuScenes. As depicted in the Figure 3(a), Weather-KITTI contains 139 million, 383 million, and 89 million LiDAR points for snow, fog, and rain noise, respectively. In comparison, as shown in Figure 3(b), Weather-NuScenes has 23 million, 112 million, and 6 million LiDAR points for snow, fog, and rain noise, respectively. This highlights the large scale and diverse weather conditions captured in our proposed datasets.
IV Denoising Network
@@ IV-A Problem Statement
IV-B The Overview of Network
In this section, we propose the TripleMixer, a robust denoising model designed for removing 3D point cloud noises under adverse weather conditions. The architecture of our model is depicted in Figure 4. The TripleMixer consists of three primary layers: 1) the Geometry Mixer (GMX) Layer, which mixes the spatial geometric relationships of neighboring points, thereby enabling the network to discern crucial local structural features and maintain important geometric details; 2) the Frequency Mixer (FMX) Layer, employing multi-scale wavelet analysis to achieve a more comprehensive representation of both noise and normal points in the frequency domain, thereby enhances the model’s ability to discern fine-grained characteristics; and 3) the Channel Mixer (CMX) Layer, which enhances the overall feature representation by mixing contextual information across channels, improving the model’s accuracy and adaptability in complex environments. Specifically, in the GMX layer, we first voxelize the LiDAR point cloud in 3D space to facilitate downsampling, then select K neighboring points and employ the attentive pooling operation to effectively capture local geometric patterns. For the FMX layer, we initially quantize the 3D features derived from the GMX layer along the X, Y, and Z axes into 2D grids. These grids are projected onto the YZ, XZ, and XY planes, creating the corresponding 2D grid images for triple plane projection. Subsequently, these images are processed using an adaptive wavelet module that provides a multi-resolution decomposition, efficiently enhancing the learning of multi-scale features. In the CMX layer, we mix feature information across different channels, synthesizing information across various directions and scales, thereby enriching the network with superior contextual information. Detailed discussions of these components will follow in subsequent sections.
@@ IV-D Frequency Mixer (FMX) Layer
Triple Plane Projection (TPP) Block . It’s aims to obtain the 2D features $x_{i}$ of the triple projection planes from the 3D features $F_{i}$ outputted by the GMX layer. To address the intensive computational demands of 3D convolution operations and the loss of height-related details in BEV projection, we introduce a novel approach by projecting the 3D point cloud along the X, Y, and Z axes onto three distinct 2D planes at varying resolutions, as illustrated in Figure 6. The TPP block utilizes sparse tensor operations to efficiently handle data. It starts by identifying non-zero indices in the point cloud to create sparse projection matrices for quantizing 3D points into 2D planes. The value of each pixel in the 2D plane represents the average of the 3D features falling within the same pixel. This triple plane projection method reduces redundancy and computational load while capturing detailed object features and enhancing spatial information, leading to more precise and accurate object characteristics.
@@
[…]
@@ IV-F Loss Function
To train our proposed model and ensure that the captured information is relevant to the point cloud denoising task, the overall loss function consists of three components: the cross-entropy loss, the lovász loss[48] and the wavelet regularization terms. These components are structured as:
$\mathcal{L}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\text{lovász }}+\mathcal{L}_{\mathrm{wr}}$ (23)
@@
where $\mathcal{L}_{\mathrm{ce}}$ represents the cross-entropy loss, assessing the accuracy of noise predictions, and $\mathcal{L}_{\text{lovász}}$ denotes the lovász loss, which is designed for improving performance in tasks with imbalanced classes. $\mathcal{L}_{\mathrm{wr}}$ represents the wavelet regularization terms, which include two term. The first wavelet regularization term minimizes the sum of the $l_{2}$ norm of the detail components across all levels, enhancing detail preservation. The second wavelet term minimizes the $l_{2}$ norm of the difference between the approximation components of consecutive levels across all levels, crucial for preserving the mean of the input signal and ensuring a proper wavelet decomposition enforcing a consistent wavelet decomposition structure, which is essential for effectively capturing multi-scale geometric information. $\lambda_{1}$ and $\lambda_{2}$ are the wavelet regularization parameters used to adjust the strength of the regularization terms. $\mathcal{D}_{t}$ represents the mean of the detail sub-bands at decomposition level $t$, while $\mathcal{A}_{t}$ and $\mathcal{A}_{t-1}$ denote the mean of the approximation sub-bands at levels $t$ and level ${t-1}$ of the lifting wavelet, respectively.
V Experiments
In this section, we first introduce the experimental datasets and their division into training, validation, and testing subsets for model training. Next, we describe the implementation details and evaluation metrics of the experiments. Moreover, we conduct extensive performance comparison experiments to benchmark our model against state-of-the-art models. Finally, comprehensive ablation studies are performed to validate the contribution of each component of our model.
V-A Datasets
We conduct qualitative tests on three large-scale LiDAR datasets: WADS[17], as well as the Weather-KITTI and Weather-NuScenes datasets introduced in this paper. These datasets contain a variety of adverse weather conditions. The details of these datasets will be introduced as follows.
WADS[17] contains LiDAR scans from 20 different sequences, with around 1300 LiDAR frames. The majority of these LiDAR scans are collected in snowy urban driving scenarios. Points in a LiDAR scan has been labeled into one of 22 semantic classes. The default labels for snow points are mapped as 110. We divided the WADS dataset [^1] The training set is: [14,15,18,20,24,28,34,36,37]. The val set is: [11,16]. The test set is: [12,13,17,22,23,26,30,35,76]. into training, validation, and testing sets according to the protocol outlined in[22]. To ensure a fair comparison for the point cloud denoising task under adverse weather conditions, we re-trained state-of-the-art models using our splits.
Table III: Splitting of Training, Validation, and Testing Sets for Our Proposed Weather-KITTI and
Weather-NuScenes Datasets.
@@ V-A Datasets
Rain Scenarios Train: [01,02,03,04,06,
@@
Weather-KITTI and Weather-NuScenes contain three different weather scenarios: snow, fog, and rain. Specifically, the Weather-KITTI dataset consists of 22 sequences for each weather scenario, while the Weather-NuScenes dataset includes 10 sequences per scenario. For model training, we divided the training, validation, and testing sets as detailed in Table V-A.
V-B Implement Details
In our experiments, we downsampled the point clouds by retaining only one point per 10 cm voxel. The resolutions for the 2D projections along the X, Y, and Z axes are set to [256, 256, 32], respectively. We optimized the model parameters using the AdamW optimizer[49] for 20 epochs with a weight decay of 0.005. The learning rate scheduler first performed a linear warmup over the initial 2 epochs, gradually increasing the learning rate from 0 to 0.001. After the warmup phase, it transitioned into a cosine annealing phase, where the learning rate decayed following a cosine curve, eventually reaching a minimum of $10^{-5}$ by the end of the last epoch. We set the batch size to 4 for each GPU and employed various data augmentation techniques to enhance model robustness, including random rotations, scaling, and random flipping along the XY plane. Additionally, we used stochastic depth with a layer drop probability of 0.2. The loss is the sum of the cross-entropy loss, the lovász loss and the wavelet regularization terms. The regularization parameters $\lambda_{1}$ and $\lambda_{2}$ are set to 0.1 for all the experiments. All results are produced on an Ubuntu 18.04 system equipped with two Intel® Xeon® Platinum 8280 CPUs, two Nvidia V100 GPUs, and 32 GB of RAM.
V-C Evaluation Metrics
Following prior work[22, 21, 20], we evaluate the point cloud denoising task using $precision=\frac{TP}{{TP+FP}}$, $recall=\frac{TP}{{TP+FN}}$, F1 score, and mean Intersection over Union (mIoU) as quantitative metrics. Specifically, the F1 score and mIoU are defined as follows:
@@ V-C Evaluation Metrics
Where $TP$ refers to the correctly identified noise points, $FP$ refers to the non-noise points incorrectly identified as noise, and $FN$ refers to the noise points that were not correctly identified.
V-D Quantitative Result
Results On WADS: We compare our model with the current state-of-the-art models, including both statistical and deep learning approaches. Specifically, SOR[16], ROR[16], DSOR[17], and DROR[30] are statistical methods. SalsaNext[32] and Cylinder3D[33] are general point cloud segmentation methods that we have adapted for binary classification to enable comparison. WeatherNet[20], 4DenoiseNet[21], and 3D-OutDet[22] are specialized point cloud denoising networks.
Table LABEL:wads presents a comprehensive comparison of denoising performance on the WADS[17] dataset, evaluating precision, recall, F1 score, and mIoU. The WADS[17] dataset represents real-world conditions, making it a robust benchmark for performance assessment. As illustrated in Table LABEL:wads, while Cylinder3D[33] achieves the highest precision with a score of 97.12, our model closely follows with a minimal difference of approximately 0.06. More importantly, our model outperforms all existing methods in other critical metrics, achieving the highest recall, F1 score, and mIoU with scores of 93.93, 94.97, and 90.73, respectively. These results underscore the superior efficacy of our model in point cloud denoising, particularly in complex real-world applications.
Table IV: Denoising Performance Results on WADS Datasets. The Proposed TripleMixer Achieves the Best mIoU.
@@ V-D Quantitative Result
Table V: Denoising Performance Results On The Proposed Weather-KITTI Datasets. Average mIoU Represents the Mean mIOU Across Three Weather Scenarios: Snow, Fog, and Rain.
@@
Table VI: Denoising Performance Results On The Proposed Weather-NuScenes Datasets. Average mIoU Represents the Mean mIOU Across Three Weather Scenarios: Snow, Fog, and Rain.
@@
Figure 7: Visualization of Denoising Performance Results on Weather-KITTI and Weather-NuScenes Datasets. Our TripleMixer demonstrates consistently stable performance across different scenarios.
Figure 8: Qualitative results among the proposed TripleMixer and state-of-the-art denoising models[21, 22, 33] on the WADS[17], the proposed Weather-KITTI and Weather-NuScenes datasets. To highlight the differences, the correct / incorrect predictions are painted in gray / red, respectively. Best viewed in colors and zoomed-in for details.
Results On Weather-KITTI: In the previous section, we demonstrated the superior performance of our denoising model on the real-world WADS[17] dataset. To further validate the good generalizability of our model, we conducted additional denoising comparison experiments on our proposed Weather-KITTI and Weather-NuScenes datasets. For a thorough evaluation, we selected the most recent and best-performing models from Table LABEL:wads for further comparative analysis.
As shown in Table LABEL:Weather-KITTI, we compared our model with six state-of-the-art methods across three scenarios in the Weather-KITTI dataset: Snow Scenarios, Fog Scenarios and Rain Scenarios. Our model achieved the highest average mIoU of 96.31 across these three scenarios, surpassing popular general semantic segmentation methods such as Cylinder3D[33], as well as the latest denoising models like 4DenoiseNet[21] and 3D-OutDet[22]. Specifically, our model’s average mIoU on the Weather-KITTI dataset is 2.2% higher than Cylinder3D[33], 16.2% higher than 4DenoiseNet[21], and 5.9% higher than 3D-OutDet[22]. It is noteworthy that our model consistently achieved the highest mIoUs of 96.46, 95.61, and 96.87 across Snow, Fog, and Rain scenarios, respectively, while DROR[30], Cylinder3D[33], and SalsaNext[32] achieved the highest Recall or Precision in the Snow, Fog, and Rain scenarios, respectively. This demonstrates the robustness of our model in point cloud denoising on the Weather-KITTI dataset.

### v2

@@ III Weather‑KITTI and Weather‑NuScenes Datasets
III-A Dataset Statistics
Figure 1 shows the semantic distributions of a single LiDAR frame captured under three different weather conditions from our Weather‑KITTI dataset, along with the corresponding clean reference frame. Weather‑induced noise points are predominantly concentrated near the LiDAR sensor, with snow producing more dispersed and irregular noise patterns than rain and fog. Table II presents the severity levels of various weather-induced corruptions configured in this study. Based on Table II, we simulated different severity levels (light, moderate, and heavy) of snow, fog, and rain. As weather severity increases, the degree of LiDAR point cloud corruption escalates accordingly. All LiDAR frames in the Weather‑KITTI and Weather‑NuScenes datasets were randomly synthesized across the three severity levels, with a comparable number of frames maintained for each level to ensure balanced data distribution. As illustrated in Figure 2(a), Weather-KITTI contains approximately 14K frames per weather type at each severity level, while Weather-NuScenes includes around 9K frames per category and severity level, as shown in Figure 2(b). In Weather‑KITTI, the proportions of degraded points at the light, moderate, and heavy severity levels across all weather classes (snow, fog, rain) are 17.55%, 30.85%, and 51.60%, respectively; In Weather‑NuScenes, the corresponding values are 14.51%, 37.21%, and 48.28%.
Weather‑KITTI, based on SemanticKITTI [36], comprises 22 sequences with 130,656 LiDAR scans. Each point cloud is segmented into 21 semantic classes, with the weather-related classes for snow, fog, and rain mapped to labels 110, 111, and 112, respectively. As SemanticKITTI provides complete semantic labels only for the first ten sequences, we restrict the use of standard semantic labels accordingly. However, our dataset provides weather class annotations across all 22 sequences. Weather-NuScenes is derived from nuScenes-lidarseg [37] and includes 84,390 LiDAR scans evenly distributed across ten sequences. It covers 19 semantic classes, along with added labels for snow, fog, and rain using the same indices as in Weather-KITTI. The top row of Figure 3 shows the total number of LiDAR points per semantic class for all sequences in our datasets: Weather-KITTI contains 139 million, 383 million, and 89 million points corrupted by snow, fog, and rain noise, respectively. Similarly, Weather-NuScenes includes 23 million, 112 million, and 6 million corrupted points under the same conditions. The bottom row of Figure 3 presents the proportion of points affected by each weather artifacts per semantic class. Weather‑KITTI exhibits more pronounced weather effects on traffic‑related categories such as cars, trucks, and motorcycles, which are typically located near the LiDAR sensor and in central road regions where particles cause stronger occlusion, scattering, and reflection. In contrast, Weather‑NuScenes shows a more balanced distribution of weather corruption across classes, reflecting its diverse urban scenes and multiple LiDAR viewpoints.
((a))
((b))
Figure 4: Spatial Distribution Characteristics of Noise Points in WADS and Our Dataset.
Table III: Intensity Distribution Characteristics of Noise Points in WADS and Our Dataset.
Intensity Range WADS (%) Ours (%)
$\left[0.0,\ 0.3\right]$ 96.63 91.92
$\left[0.3,\ 0.6\right]$ 3.01 7.54
$\left[0.6,\ 1.0\right]$ 0.36 0.54
III-B Dataset Comparison
To assess the practical reliability of our simulated datasets, we compare them with WADS [18], which provides real‑world LiDAR sequences collected in snowy conditions with annotated noise labels. We randomly sample 1,000 scans from each dataset and analyze two key noise properties: spatial and intensity distributions. As illustrated in Figure 4, both datasets exhibit similar spatial characteristics. Approximately 90% of noise points are concentrated within a range of 3 to 10 meters from the LiDAR sensor, with a gradual decline beyond 10 meters and negligible presence past 20 meters. This consistency indicates that our simulations effectively reproduce real‑world occlusion and backscatter effects. Moreover, the intensity distribution of noise points, summarized in Table III, further demonstrates the similarity. Over 90% of noise points in both datasets fall within the low-intensity range [0.0, 0.3], indicating consistent degradation characteristics. Minor differences in the mid-range interval [0.3, 0.6] may be attributed to variations in LiDAR hardware or scene-specific factors, yet the overall distribution trend remains well aligned. These findings further validate that our synthetic weather corruption effectively replicates real-world noise patterns in both spatial and intesity properties.
@@ III-B Dataset Comparison
Figure 5: Overview of the proposed TripleMixer denoising network, which consists of three key layers: the Geometry Mixer (GMX), Frequency Mixer (FMX), and Channel Mixer (CMX). TripleMixer is trained solely using point-wise weather labels.
IV TripleMixer Model
@@ IV-A Problem Statement
IV-B Overall Framework
In this section, we propose the TripleMixer, a robust denoising model designed for removing 3D point cloud noises under adverse weather conditions. Its architecture is illustrated in Figure 5 and comprises three primary layers: 1) the Geometry Mixer (GMX) Layer, which mixes the spatial geometric relationships of neighboring points, thereby enabling the network to discern crucial local structural features and maintain important geometric details; 2) the Frequency Mixer (FMX) Layer, employing multi-scale wavelet analysis to achieve a more comprehensive representation of both noise and normal points in the frequency domain; and 3) the Channel Mixer (CMX) Layer, which enhances the overall feature representation by mixing contextual information across channels, improving the model’s accuracy and adaptability in complex environments. Specifically, in the GMX layer, we first voxelize the LiDAR point cloud in 3D space to facilitate downsampling, then select K neighboring points and employ the attentive pooling operation to effectively capture local geometric patterns. For the FMX layer, we initially quantize the 3D features derived from the GMX layer along the X, Y, and Z axes into 2D grids. These grids are projected onto the YZ, XZ, and XY planes, creating the corresponding 2D grid images for triple plane projection. Subsequently, these images are processed using an adaptive wavelet module that provides a multi-resolution decomposition, efficiently enhancing the learning of multi-scale features. In the CMX layer, we mix feature information across different channels, synthesizing information across various directions and scales, thereby enriching the network with superior contextual information. Detailed discussions of these components will follow in subsequent sections.
@@ IV-D Frequency Mixer (FMX) Layer
Triple Plane Projection (TPP) Block. It aims to obtain the 2D features $x_{i}$ of the triple projection planes from the 3D features $F_{i}$ outputted by the GMX layer. To address reduce the computational complexity of 3D convolution operations and the loss of height-related details in BEV projection, we introduce a novel approach by projecting the 3D point cloud along the X, Y, and Z axes onto three distinct 2D planes at varying resolutions, as illustrated in Figure 7. The TPP block utilizes sparse tensor operations to efficiently handle data. It starts by identifying non-zero indices in the point cloud to create sparse projection matrices for quantizing 3D points into 2D planes. The value of each pixel in the 2D plane represents the average of the 3D features falling within the same pixel. This triple plane projection method reduces redundancy and computational load while capturing detailed object features and enhancing spatial information, leading to more precise and accurate object characteristics.
@@
$x^{t}_{HL},x^{t}_{HH}]=\mathit{VWD}\left(d^{t}\right),$ (8)
$x^{t}_{LL},x^{t}_{LH}]=\mathit{VWD}\left(c^{t}\right),$ (9)
@@ IV-F Loss Function
To train our proposed model and ensure that the captured information is relevant to the point cloud denoising task, the overall loss function consists of three components: the cross-entropy loss, the lovasz loss[42] and the wavelet regularization terms. These components are structured as:
$\mathcal{L}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\text{lovasz}}+\mathcal{L}_{\mathrm{wr}}$ (13)
@@
where $\mathcal{L}_{\mathrm{ce}}$ represents the cross-entropy loss, assessing the accuracy of noise predictions, and $\mathcal{L}_{\text{lovasz}}$ denotes the lovasz loss, which is designed for improving performance in tasks with imbalanced classes. $\mathcal{L}_{\mathrm{wr}}$ represents the wavelet regularization terms, which include two term. The first wavelet regularization term minimizes the sum of the $l_{2}$ norm of the detail components across all levels, enhancing detail preservation. The second wavelet term minimizes the $l_{2}$ norm of the difference between the approximation components of consecutive levels across all levels, crucial for preserving the mean of the input signal and ensuring a proper wavelet decomposition enforcing a consistent wavelet decomposition structure, which is essential for effectively capturing multi-scale geometric information. $\lambda_{1}$ and $\lambda_{2}$ are the wavelet regularization parameters used to adjust the strength of the regularization terms. $\mathcal{D}_{t}$ represents the mean of the detail sub-bands at decomposition level $t$, while $\mathcal{A}_{t}$ and $\mathcal{A}_{t-1}$ denote the mean of the approximation sub-bands at levels $t$ and level t1{t-1} of the lifting wavelet, respectively.
V Denoising and Perception Benchmarks under Adverse Weather
In this section, we establish a Denoising benchmark to evaluate the performance of our denoising model and introduce three downstream LiDAR perception benchmarks: Semantic Segmentation (SS), Place Recognition (PR), and Object Detection (OD), to assess the generalization of state‑of‑the‑art perception models under adverse weather and the effectiveness of our denoising model as a preprocessing step. Notably, in all downstream benchmarks, our denoising model is trained in a supervised manner solely on our Weather‑KITTI and Weather‑NuScenes datasets using only point‑wise weather labels. Meanwhile, all perception models are directly tested on real‑world adverse‑weather datasets without any retraining or fine‑tuning. This design highlights the strong transferability of our datasets and the practical value of the proposed denoising models for real‑world applications. All experiments were conducted on an Ubuntu 18.04 with Intel Xeon Platinum 8280 CPUs, two Nvidia V100 GPUs, and 32 GB RAM.
Table IV: Splitting of Training, Validation, and Testing Sets for Our Datasets.
@@ V Denoising and Perception Benchmarks under Adverse Weather
Rain Scenarios Train:
@@
V-A Denoising Benchmark
Benchmark Setting: We conduct denoising experiments on the real‑world extreme‑weather dataset WADS [18] and our Weather‑KITTI and Weather‑NuScenes datasets. WADS[18] comprises 1,300 LiDAR scans across 20 sequences, collected in snowy urban driving scenarios. We divided the WADS dataset [^1] The training set is: [14,15,18,20,24,28,34,36,37]. The val set is: [11,16]. The test set is: [12,13,17,22,23,26,30,35,76]. into training, validation, and testing sets according to the protocol outlined in[14]. To ensure a fair comparison, we re-trained state-of-the-art models using our splits. For Weather‑KITTI and Weather‑NuScenes, the dataset partitions for training, validation, and testing are provided in Table V. During training, point clouds were downsampled to a 10 cm voxel grid and projected onto the X, Y, and Z planes at [256, 256, 32] resolutions. Models were trained for 30 epochs using AdamW [43] with a weight decay of 0.005, a batch size of 4 per GPU, and stochastic depth (drop probability 0.2). The learning rate warmed up linearly to 0.001 over 2 epochs, then decayed to $10^{-5}$ with cosine annealing. The loss combined cross‑entropy, Lovasz, and wavelet regularization with $\lambda_{1}=\lambda_{2}=0.1$.
Following prior work [14, 21, 20], we evaluate the point cloud denoising task using precision, recall, F1 score, and mean Intersection‑over‑Union (mIoU) as quantitative metrics, where $precision=\frac{TP}{TP+FP}$ and $recall=\frac{TP}{TP+FN}$. Specifically, the F1 score and mIoU are defined as
@@ V-A Denoising Benchmark
where $TP$ denotes correctly identified noise points, $FP$ denotes non‑noise points incorrectly identified as noise, and $FN$ denotes noise points that were not correctly identified.
Results On WADS: We compare our model with the current state‑of‑the‑art models, with a primary focus on point cloud denoising networks. Specifically, WeatherNet [20], 4DenoiseNet [21], and 3D‑OutDet [14] are specialized point cloud denoising networks. In addition, SOR [17], ROR [17], DSOR [18], and DROR [25] are statistical methods, while SalsaNext [44], Cylinder3D [45], and LSK3DNet [46] are point cloud segmentation methods that we have adapted for binary classification to enable comparison. Table LABEL:wads presents a comprehensive comparison of denoising performance on the WADS[18] dataset, evaluating precision, recall, F1 score, and mIoU. As illustrated in Table LABEL:wads, while LSK3DNet[46] achieves the highest precision with a score of 97.16, our model surpasses all existing methods in other critical metrics, achieving the highest recall, F1 score, and mIoU with scores of 93.93, 95.13, and 90.73, respectively. These results underscore the superior efficacy of our model in point cloud denoising, particularly in complex real-world applications.
Table V: Denoising Performance Results on WADS Datasets.
@@
Figure 8: Qualitative results of the Denoising benchmark on the WADS [18], Weather‑KITTI (WK), and Weather‑NuScenes (WN) datasets.
Semantic point colors follow the scheme in Figure 1, with red points indicating weather‑induced noise.
Table VI: Denoising Performance On The Proposed Weather-KITTI Datasets. Average mIoU Represents the Mean mIOU Across Three Weather Scenarios.
@@
Table VII: Denoising Performance Results On The Proposed Weather-NuScenes Datasets. Average mIoU Represents the Mean mIOU Across Three Weather Scenarios: Snow, Fog, and Rain.
@@
Results On Weather-KITTI: To further validate the good generalizability of our model, we conducted additional denoising comparison experiments on our proposed Weather-KITTI and Weather-NuScenes datasets. For a thorough evaluation, we selected the most recent and best-performing models from Table LABEL:wads for further comparative analysis. As shown in Table LABEL:Weather-KITTI, we compared our model with seven state-of-the-art methods across three scenarios in the Weather-KITTI dataset: Snow Scenarios, Fog Scenarios and Rain Scenarios. Our model achieved the highest average mIoU of 96.31, outperforming general semantic segmentation methods such as LSK3DNet [46] and advanced denoising models like 4DenoiseNet [21] and 3D‑OutDet [14]. Specifically, our model’s average mIoU exceeded LSK3DNet, 4DenoiseNet, and 3D‑OutDet by 1.7%, 16.2%, and 5.9%, respectively. Notably, our model achieved the highest mIoUs of 96.46, 95.61, and 96.87 in Snow, Fog, and Rain scenarios, respectively, whereas Cylinder3D [45], and SalsaNext [44] attained the highest Precision in individual scenarios. This likely reflects their conservative filtering or task‑specific feature modeling, which improves precision by suppressing false positives but often reduces recall. In contrast, our Frequency Mixer (FMX) module performs multi‑scale frequency decomposition through a learnable lifting‑wavelet design, enabling the selective retention of task‑relevant high‑frequency details. This design prioritizes structural completeness over aggressive filtering, thereby striking a better balance between precision and recall and ensuring robust and consistent denoising across diverse weather conditions.